OpenStudy (anonymous):

Verify the identity. tan (x+pi/2) = -cot x

OpenStudy (anonymous):

tan(x+pi/2) = (sin(x+pi/2)/(soc(x+pi/2) right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ummmmm ahaaa sin(pi/2 + x) = - cos (x) and cos (pi/2+x) = sin(x)

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

i give you a picture hint

OpenStudy (anonymous):